OpenStudy (anonymous):

Confirm that f and g are inverses by showing that f(g(x)) = x and g(f(x)) = x. f(x) = the quantity x minus seven divided by the quantity x plus three. and g(x) = quantity negative three x minus seven divided by quantity x minus one.

OpenStudy (anonymous):

\[f(x)=\frac{ x-7 }{ x+3 } g(x)=\frac{ -3x-7 }{ x-1 }\]

OpenStudy (anonymous):

prepare to do a raft of algebra

OpenStudy (anonymous):

ah okay thank you for your help in advance btw :)

OpenStudy (anonymous):

\[f(x)=\frac{ x-7 }{ x+3 } \] and \[g(x)=\frac{ -3x-7 }{ x-1 }\] right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

this is amazingly easy for me to write the composition one way by using copy and paste i am going to past \(g(x)=\frac{ -3x-7 }{ x-1 }\)where every i see an \(x\) in \(f(x)=\frac{ x-7 }{ x+3 }\)

OpenStudy (anonymous):

\[f(x)=\frac{ x-7 }{ x+3 }\] \[f(g(x))=\frac{ \frac{ -3x-7 }{ x-1 }-7 }{ \frac{ -3x-7 }{ x-1 }+3 }\]

OpenStudy (anonymous):

okay i did that and now do i just solve?

OpenStudy (anonymous):

now comes the raft of algebra part

OpenStudy (anonymous):

which will end up with an orgy of cancellation since your answer should be \(x\)

OpenStudy (anonymous):

it is not really that bad multiply top and bottom of that complex fraction by \(x-1\) to clear the denominator don't forget the distributive law

OpenStudy (anonymous):

okay let me just do it real quick

OpenStudy (anonymous):

would i do the same for the g(f(x)) = x. later on?

OpenStudy (anonymous):

yes, just clear the denominators then all will go bye bye

OpenStudy (anonymous):

ah thank you so much!

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

wait im so lost what happens with the -7 and 3?

OpenStudy (anonymous):

don't forget the distributive law !!

OpenStudy (anonymous):

????

OpenStudy (anonymous):

\[\frac{ \frac{ -3x-7 }{ x-1 }-7 }{ \frac{ -3x-7 }{ x-1 }+3 }\times \frac{x-1}{x-1}\] \[=\frac{-3x-7(x-1)}{-3x-7+3(x-1)}\] is a start

OpenStudy (anonymous):

the distributive law in action denominators go, but you still have to distribute now distribute again, the -7 up top and the 3 below then combine like terms

OpenStudy (anonymous):

okay at the denominator should i add the 7 and 3 before distrubuting or no?

OpenStudy (anonymous):

distribute first

OpenStudy (anonymous):

numerator should be \(-10x+7\) if you do it correctly

OpenStudy (anonymous):

okay just a min

OpenStudy (anonymous):

the denom i got -10

OpenStudy (anonymous):

yeah that is right, and i lied, the numerator is \(-3x-7-7(x-1)=-3x-7-7x+7=-10x\)

OpenStudy (anonymous):

and of course \(\frac{-10x}{-10}=x\) as needed

OpenStudy (anonymous):

ah okay thats easy in reality thanks! ill do the other side now

OpenStudy (anonymous):

good luck it is real similar

OpenStudy (anonymous):

i just am stuck a little after muliplying x+3/x+3

OpenStudy (anonymous):

\[g(x)=\frac{ -3x-7 }{ x-1 }\] \[g(f(x))=\frac{ -3\frac{ x-7 }{ x+3 }-7 }{ \frac{ x-7 }{ x+3 }-1 }\]

OpenStudy (anonymous):

yeah i did that already

OpenStudy (anonymous):

multiply by \(x-3\) what did you get in the numerator ?

OpenStudy (anonymous):

why x-3?

OpenStudy (anonymous):

typo i meant \(x+3\)

OpenStudy (anonymous):

okay making sure ahha

OpenStudy (anonymous):

-3(x-7)-7 is my numerator?

OpenStudy (anonymous):

before distributing etc you should be looking at \[-3(x-7)-7(x+3)\]

OpenStudy (anonymous):

forgot that distributive law didn't you?

OpenStudy (anonymous):

ah yes

OpenStudy (anonymous):

it is the distributive LAW, not the distributive option

OpenStudy (anonymous):

hahaha yes so my numerator is:(-3x+21)-7x-21

OpenStudy (anonymous):

numerator comes to -10x

OpenStudy (anonymous):

got it perfect! thanks!

OpenStudy (anonymous):

yw